OpenStudy (anonymous):

Which expression is the simplified form of √49a^4

OpenStudy (saifoo.khan):

\[\large \sqrt{49} \times \sqrt{a^4}\]

OpenStudy (anonymous):

So?

OpenStudy (saifoo.khan):

so, solve it.

OpenStudy (anonymous):

Lol you're funny haha

OpenStudy (saifoo.khan):

LOL. ^_~